OpenStudy (anonymous):

y=3x^2-12x-15 Express in vertex form: 3(x-h)^2+k

sam (.sam.):

Do you know completing the square?

OpenStudy (anonymous):

a=3, b=-12, c=-15 h=-b/(2a) k=3(h)^2-12(h)-15

OpenStudy (anonymous):

the formula for h is to get the x coordinate of the vertex. then to get k, just plug that h into the original equation.

OpenStudy (anonymous):

ok thank you!